# Spectral Clustering with Likelihood Refinement is Optimal for Latent Class Recovery

### v1

@@ Abstract
Latent class models are widely used for identifying unobserved subgroups from multivariate categorical data in social sciences, with binary data as a particularly popular example. However, accurately recovering individual latent class memberships and determining the number of classes remains challenging, especially when handling large-scale datasets with many items. This paper proposes a novel two-stage algorithm for latent class models with high-dimensional binary responses. Our method first initializes latent class assignments by an easy-to-implement spectral clustering algorithm, and then refines these assignments with a one-step likelihood-based update. This approach combines the computational efficiency of spectral clustering with the improved statistical accuracy of likelihood-based estimation. We establish theoretical guarantees showing that this method leads to optimal latent class recovery and exact clustering of subjects under mild conditions. Additionally, we propose a simple consistent estimator for the number of latent classes. Extensive experiments on both simulated data and real data validate our theoretical results and demonstrate our method’s superior performance over alternative methods.
Keywords: High-dimensional data; Joint maximum likelihood estimation; Latent class model; Spectral clustering;
@@ 1 Introduction
Latent class models (LCMs, Lazarsfeld,, 1968; Goodman,, 1974) are popular tools for uncovering unobserved subgroups from multivariate categorical responses, with broad applications in social sciences (Korpershoek et al.,, 2015; Berzofsky et al.,, 2014; Wang and Hanges,, 2011; Zeng et al.,, 2023). At the core of an LCM, we aim to identify $K$ latent classes of $N$ individuals based on their categorical responses to $J$ items. In this paper, we focus on the widely used LCMs with binary responses, which are prevailing in educational assessments (correct/wrong responses to test questions) and social science surveys (yes/no responses to survey items). Accurately and efficiently recovering latent classes in large $N$ and large $J$ settings remain a significant statistical and computational challenge.
@@
Traditional approaches to LCM are typically based on maximum likelihood estimation. In the classical random-effect formulation, the latent class labels of subjects are treated as random variables and the model is estimated by the marginal maximum likelihood (MML) approach, in which the latent labels are marginalized out. The expectation-maximization (EM) algorithm (Dempster et al.,, 1977) is often used for this purpose but becomes computationally expensive and lacks theoretical guarantees in the modern regime where both $N$ and $J$ are large. In contrast, the joint maximum likelihood (JML) approach adopts a fixed-effect perspective, treating the latent class labels as unknown parameters to be estimated. However, due to the inherent non-convexity of the joint likelihood function, JML is still highly sensitive to initialization and prone to be trapped in local optima. Recently, Zeng et al., (2023) showed that JML is consistent under the large $N$ and large $J$ regime. Nonetheless, they do not provide any algorithmic guarantee, and the convergence rate established therein is polynomial in $J$, whose optimality remains unclear.
@@
In this paper, we consider a different approach related to spectral clustering, which leverages the intrinsic approximate low-rank structure of the data matrix and is computationally efficient. In psychometrics and related fields, the common heuristic practice for spectral clustering is to first construct a similarity matrix among individuals or items, apply eigenvalue decomposition to it, and then perform K-means clustering of the eigenspace embeddings (Von Luxburg,, 2007; Chen et al.,, 2017). More recently, Chen and Li, (2019) directly performs the singular value decomposition (SVD) on a regularized version of the data matrix itself. Similar ideas have also been developed for nonlinear item factor analysis (Zhang et al.,, 2020) and the grade-of-membership model (Chen and Gu,, 2024).
[…]
@@
This hybrid approach simultaneously takes advantage of the scalability of spectral methods and leverages the Bernoulli likelihood information to achieve a vanishing exponential error rate for clustering. We provide rigorous theoretical guarantees for our method. In summary, our method is both computationally fast and statistically optimal in the regime with a large number of items. Therefore, this work provides a useful tool to aid psychometric researchers and practitioners to perform latent class analysis of modern high-dimensional data.
The rest of this article is organized as follows. After reviewing the setup of LCMs for binary responses in Section 2, we discuss the performance of spectral clustering and propose our main algorithm in Section 3. In Section 4, we establish theoretical guarantees and statistical optimality of our algorithm. We evaluate our method’s performance in numerical experiments in Section 5.
@@
[…]
@@ 2 Latent Class Model with Binary Responses
A standard approach for maximizing the marginal likelihood is the EM algorithm (Dempster et al.,, 1977). However, EM is highly sensitive to initialization. Poor initial values often lead to convergence at local maxima, so practitioners resort to multiple random restarts at substantial computational cost. Moreover, when both $J$ and $K$ grow large, the computation for evaluating the objective in (2) becomes numerically unstable due to its log-sum-of-products structure.
@@
The simple structure of (3) implies that the optimization of JML can be decoupled across $[N]$ and $[J]$. While JML is known to be inconsistent in low-dimensional settings (Neyman and Scott,, 1948), the recent study Zeng et al., (2023) has shown that JML can achieve consistent estimation of the latent labels in high-dimensional regime. This property makes the JML approach appealing for large-scale modern datasets.
@@
where ${\mathfrak{S}}_{K}$ denotes the set of all permutation maps $\pi$ from $[K]\to[K]$. This metric accounts for the inherent symmetry in clustering structure.
@@ 3.1 Spectral clustering for the latent class model
Thus, data points in the same latent class have identical row vectors in $\mathbf{U}\mathbf{\Sigma}$, namely one of the $K$ distinct vectors $\left\{\mathbf{\Theta}_{:,k}^{\top}\mathbf{V}\right\}_{k=1}^{K}$. Hence the true latent class label information can be readily read off from rows of $\mathbf{U}\mathbf{\Sigma}$. In the realistic noisy data setting, denote by ^$\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}\widehat{}\mathbf{V}^{\top}$ the top-$K$ SVD of the data matrix $\mathbf{R}$. Then intuitively, $\mathbf{U}\mathbf{\Sigma}$ would be well-approximated by ^$\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ provided that $\mathbf{E}$ is relatively small in magnitude, and the row vectors of ^$\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ are noisy point clouds distributed around the $K$ points. Therefore, K-means clustering can deliver reasonable estimates of class labels. We illustrate this point using an numerical example in Figure 1, where the row vectors of $\mathbf{U}\mathbf{\Sigma}$ (left) and ^$\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ (right) are plotted as points in $\mathbb{R}^{2}$, respectively.
The above high-level idea can be formalized as the spectral clustering algorithm (Algorithm 1), investigated in Zhang and Zhou, (2024) for sub-Gaussian mixture models. Algorithm 1 first obtains the rank-$K$ SVD of $\mathbf{R}$ denoted by ^$\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}\widehat{}\mathbf{V}^{\top}$, and then performs K-means clustering on the rows of ^$\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$, the left singular vectors weighted by the singular values.
Figure 1: An illustration for spectral clustering: row vectors of $\mathbf{U}\mathbf{\Sigma}$ (left) and ^$\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ (right). Setting: $N=500$, $J=250$, $K=3$.
@@
2Perform rank-$K$ SVD on $\mathbf{R}$ to obtain $\sum_{i=1}^{K}\widehat{\lambda}_{i}\widehat{\bm{u}}_{i}\widehat{\bm{v}}_{i}^{\top}$, where $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq\cdots\geq\widehat{\lambda}_{K}\geq 0$, ^$\widehat{}\mathbf{U}=\left({\bm{u}}_{1},\cdots,{\bm{u}}_{K}\right)\in\mathbb{R}^{N\times K}$ and ^$\widehat{}\mathbf{V}=\left({\bm{v}}_{1},\cdots,{\bm{v}}_{K}\right)\in\mathbb{R}^{J\times K}$.
Perform $K$-means on the rows of ^$\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ (or equivalently, $\mathbf{R}\widehat{}\mathbf{V}$), i.e.,
@@ 3.2 One-step Likelihood Refinement with Spectral Initialization
    (Assignment) Given ^$\widehat{}\mathbf{\Theta}$, the objective function $L({\bm{s}},\widehat{}\mathbf{\Theta}\mid\mathbf{R})$ can decoupled into $N$ terms each corresponding to an individual data point, and it is equivalent to assign label $k\in[K]$ to sample $i$ which maximize the log-likelihood independently for each $i\in[N]$:
@@
$\widehat{\theta}_{j,k}=\frac{\sum_{i\in[N]}R_{i,j}\mathbb{I}\left(\widetilde{s}_{i}=k\right)}{\sum_{i\in[N]}\mathbb{I}\left(\widetilde{s}_{i}=k\right)},\quad j\in[J],~{}k\in[K].$ (7)
@@
$\widehat{s}_{i}=\operatorname*{arg\,max}_{k\in[K]}\sum_{j\in[J]}\Big{[}R_{i,j}\log\Big{(}\widehat{\theta}_{j,k}\Big{)}+(1-R_{i,j})\log\Big{(}1-\widehat{\theta}_{j,k}\Big{)}\Big{]},\quad i\in[N].$ (8)
@@
Interestingly, our two-stage procedure resonates with the spirit of the “initialize-and-refine” scheme that has been used in network analysis for stochastic block models (SBMs) (Gao et al.,, 2017; Lyu et al.,, 2023; Chen et al.,, 2022), while the model considered here differs fundamentally in that LCMs handle high-dimensional response data without relying on pairwise interaction structures inherent to SBMs.
In practice, we observe that performing multiple likelihood-based refinement steps may lead to improved convergence and performance compared to the one-step refinement procedure. Thus, we also consider a multiple-step refinement variant of Algorithm 2, which we denote by SOLA+, in the later simulation studies. We set the number of “Maximization $+$ Assignment” refinement steps for SOLA+ to be $10$ in all our simulations.
@@ 3.3 Extension: Classification EM refinement
Our proposed framework can be readily extended to one-step Classification EM (CEM) refinement, where CEM is another widely used approach to estimate latent class models (Celeux and Govaert,, 1992; Zeng et al.,, 2023). It can be regarded is a variant of the standard EM algorithm designed specifically for clustering. Unlike traditional EM, which calculates “soft” assignments by estimating the probability that each observation belongs to every latent class, CEM uses “hard” assignments: each observation is directly assigned to the class with the highest posterior probability. By explicitly modeling the mixture proportion parameters for the latent classes, CEM adjusts the likelihood to account for varying cluster sizes. Concretely, in our setting, one can adapt the one-step refinement in Algorithm 2 to a CEM step by including class proportions. Besides updating $\widehat{\theta}_{j,k}$ as in (7), we also estimate the proportion parameters $\widehat{{\bm{p}}}=(\widehat{p}_{1},\ldots,\widehat{p}_{K})\in[0,1]^{K}$ by
@@
This one‐step CEM update replaces the likelihood‐refinement stage of our two‐stage estimator: start with spectral clustering, then perform the CEM step above. We provide its theoretical guarantee in Section 4.

### v2

# Spectral Clustering with Likelihood Refinement for High-dimensional Latent Class Recovery

@@ Abstract
Latent class models are widely used for identifying unobserved subgroups from multivariate categorical data in social sciences, with binary data as a particularly popular example. However, accurately recovering individual latent class memberships remains challenging, especially when handling high-dimensional datasets with many items. This work proposes a novel two-stage algorithm for latent class models suited for high-dimensional binary responses. Our method first initializes latent class assignments by an easy-to-implement spectral clustering algorithm, and then refines these assignments with a one-step likelihood-based update. This approach combines the computational efficiency of spectral clustering with the improved statistical accuracy of likelihood-based estimation. We establish theoretical guarantees showing that this method is minimax-optimal for latent class recovery in the statistical decision theory sense. The method also leads to exact clustering of subjects with high probability under mild conditions. As a byproduct, we propose a computationally efficient consistent estimator for the number of latent classes. Extensive experiments on both simulated data and real data validate our theoretical results and demonstrate our method’s superior performance over alternative methods.
Keywords: High-dimensional data; Joint maximum likelihood estimation; Latent class model; Spectral clustering.
@@ 1 Introduction
Latent class models (LCMs, Lazarsfeld, 1968; Goodman, 1974) are popular tools for uncovering unobserved subgroups from multivariate categorical responses, with broad applications in social and behavioral sciences (Korpershoek et al., 2015; Berzofsky et al., 2014; Wang and Hanges, 2011; Zeng et al., 2023). At the core of an LCM, we aim to identify $K$ latent classes of $N$ individuals based on their categorical responses to $J$ items. In this paper, we focus on the widely used LCMs with binary responses, which are prevailing in educational assessments (correct/wrong responses to test questions) and social science surveys (yes/no responses to survey items). Accurately and efficiently recovering latent classes in large $N$ and large $J$ settings remains a significant statistical and computational challenge.
@@
Traditional approaches to LCM are typically based on maximum likelihood estimation. In the classical random-effect formulation, the latent class labels of subjects are treated as random variables and the model is estimated by the marginal maximum likelihood (MML) approach, in which the latent labels are marginalized out. The expectation-maximization (EM) algorithm (Dempster et al., 1977) is often used for this purpose, but it becomes computationally expensive and lacks theoretical guarantees in the modern regime where both $N$ and $J$ are large. In contrast, the joint maximum likelihood (JML) approach adopts a fixed-effect perspective, treating the latent class labels as unknown parameters to be estimated. However, due to the inherent non-convexity of the joint likelihood function, JML is still highly sensitive to initialization and prone to be trapped in local optima. Recently, Zeng et al. (2023) showed that JML is consistent under the large $N$ and large $J$ regime. Nonetheless, they do not provide any algorithmic guarantee, and the convergence rate established therein is polynomial in $J$, whose optimality remains unclear.
@@
In this paper, we consider a different approach related to spectral clustering, which leverages the intrinsic approximate low-rank structure of the data matrix and is computationally efficient. In psychometrics and related fields, the common heuristic practice for spectral clustering is to first construct a similarity matrix among individuals or items, apply eigenvalue decomposition to it, and then perform K-means clustering of the eigenspace embeddings (Von Luxburg, 2007; Chen et al., 2017). More recently, Chen and Li (2019) directly performs the singular value decomposition (SVD) on a regularized version of the data matrix itself. Similar ideas have also been developed for nonlinear item factor analysis (Zhang et al., 2020) and the grade-of-membership model (Chen and Gu, 2024).
We next explain our idea of directly adapting the SVD to perform spectral clustering for binary-response LCMs. We collect the binary responses in an $N\times J$ matrix $\mathbf{R}$ with binary entries, then under the typical LCM assumption, the $\mathbf{R}$ can be written as as $\mathbf{R}=\mathbb{E}[\mathbf{R}]+\mathbf{E}$, where $\mathbb{E}[\mathbf{R}]$ is the low-rank “signal” matrix due to the existence of $K$ latent classes, and $\mathbf{E}$ is a mean-zero “noise” matrix. Denote by $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}\widehat{}\mathbf{V}^{\top}$ the truncated rank-$K$ SVD of $\mathbf{R}$, then $\mathbf{R}\approx\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}\widehat{}\mathbf{V}^{\top}$ approximately captures the “signal” part. In particular, we first extract $\widehat{}\mathbf{U}$, the leading $K$ left singular vectors of the data matrix to perform dimensionality reduction. Then we apply K-means clustering algorithm on the $N$ rows of $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$, the left singular vectors weighted by their corresponding singular values $\widehat{}\mathbf{\Sigma}$, to obtain the latent class label estimates. While variants of this simple yet efficient algorithm have attracted significant attention in network analysis (Zhang, 2024) and sub-Gaussian mixture models (Zhang and Zhou, 2024), it has not been widely explored or studied in the context of LCMs. The only exception is the recent work Lyu et al. (2025), which extended spectral clustering to LCMs with individual-level degree heterogeneity and showed that spectral methods can lead to exact recovery of the latent class labels in certain regime. However, whether the spectral method gives an optimal way for estimating latent class labels in LCMs remains unknown. In this paper, we resolve this question by proposing a novel two-stage method:
@@
This hybrid approach simultaneously takes advantage of the scalability of spectral methods and leverages the Bernoulli likelihood information to achieve a vanishing exponential error rate for clustering. We provide rigorous theoretical guarantees for our method. In summary, our method is both computationally fast and statistically optimal in the regime with a large number of items. Therefore, this work provides a useful tool to aid psychometric researchers and practitioners to perform latent class analysis of modern high-dimensional data. We emphasize that our method has no requirement on the magnitude of $J$ compared to $N$, and even allows $J$ to be much larger than $N$.
The rest of this article is organized as follows. After reviewing the setup of LCMs for binary responses in Section 2, we discuss the performance of spectral clustering and propose our main algorithm in Section 3. In Section 4, we establish theoretical guarantees and statistical optimality of our algorithm. We evaluate our method’s performance in numerical experiments in Section 5. The proofs of the theoretical results are included in the Appendix.
@@
We use bold capital letters such as $\mathbf{A},\mathbf{B},\cdots$ to denote matrices and bold lower-case ${\bm{a}},{\bm{b}},\cdots$ to denote vectors. For any positive integer $m$, denote $m]:=\{1,\cdots,m\}$. For any matrix $\mathbf{A}\in\mathbb{R}^{N\times J}$ and any $i\in[N]$ and $j\in[J]$, we use $A_{i,j}$ to denote its entry on the $i$-th row and $j$-th column, and use $\mathbf{A}_{i,:}$ (or $\mathbf{A}_{:,j}$) to denote its $i$-th row (or $j$-th column) vector. Let $\lambda_{k}\left(\mathbf{A}\right)$ denote the $k$-th largest singular value of $\mathbf{A}$ for $k=1,\cdots,\min\left\{N,J\right\}$. Let $\left\|\cdot\right\|$ denote the spectral norm (operator norm) for matrices and $\ell_{2}$ norm for vectors, and $\left\|\cdot\right\|_{\rm F}$ denote Frobenius norm for matrices. For two non-negative sequences $\{a_{N}\},\{b_{N}\}$, we write $a_{N}\lesssim b_{N}$ (or $a_{N}\gtrsim b_{N}$) if there exists some constant $C>0$ independent of $N$ such that $a_{N}\leq Cb_{N}$ (or $b_{N}\leq Ca_{N}$); $a_{N}\asymp b_{N}$ if $a_{N}\lesssim b_{N}$ and $b_{N}\lesssim a_{N}$ hold simultaneously; $a_{N}=o(b_{N})$ if $b_{N}>0$ and $a_{N}/b_{N}\rightarrow 0$.
@@ 2 Latent Class Model with Binary Responses
A standard approach for maximizing the marginal likelihood is the EM algorithm (Dempster et al., 1977). However, EM is highly sensitive to initialization. Poor initial values often lead to convergence at local maxima, so practitioners resort to multiple random restarts at substantial computational cost. Moreover, when both $J$ and $K$ grow large, the computation for evaluating the objective in (2) becomes numerically unstable due to its log-sum-of-products structure.
@@
The simple structure of (3) implies that the optimization of JML can be decoupled across $[N]$ and $[J]$. While JML is known to be inconsistent in low-dimensional settings (Neyman and Scott, 1948), the recent study Zeng et al. (2023) has shown that JML can achieve consistent estimation of the latent labels in the high-dimensional regime. This property makes the JML approach appealing for large-scale modern datasets. It is worth noting that JML can be regarded as a version of classification EM (Celeux and Govaert, 1992), but with the difference of omitting the class proportion parameters. See Section 3.3 for more details.
@@
where ${\mathfrak{S}}_{K}$ denotes the set of all permutation maps $\pi$ from $K]\to[K]$. This metric accounts for the inherent symmetry in clustering structure.
@@ 3.1 Spectral clustering for the latent class model
Thus, data points in the same latent class have identical row vectors in $\mathbf{U}\mathbf{\Sigma}$, namely one of the $K$ distinct vectors $\left\{\mathbf{\Theta}_{:,k}^{\top}\mathbf{V}\right\}_{k=1}^{K}$. Hence the true latent class label information can be readily read off from rows of $\mathbf{U}\mathbf{\Sigma}$. In the realistic noisy data setting, denote by $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}\widehat{}\mathbf{V}^{\top}$ the top-$K$ SVD of the data matrix $\mathbf{R}$. Then intuitively, $\mathbf{U}\mathbf{\Sigma}$ would be well-approximated by $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ provided that $\mathbf{E}$ is relatively small in magnitude, and the row vectors of $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ are noisy point clouds distributed around the $K$ points. Therefore, K-means clustering can deliver reasonable estimates of class labels. We illustrate this point using an numerical example in Figure 1, where the row vectors of $\mathbf{U}\mathbf{\Sigma}$ (left) and $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ (right) are plotted as points in $\mathbb{R}^{2}$, respectively.
The above high-level idea can be formalized as the spectral clustering algorithm (Algorithm 1), investigated in Zhang and Zhou (2024) for sub-Gaussian mixture models. Algorithm 1 first obtains the rank-$K$ SVD of $\mathbf{R}$ denoted by $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}\widehat{}\mathbf{V}^{\top}$, and then performs K-means clustering on the rows of $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$, the left singular vectors weighted by the singular values.
Remark 1.
An equivalent formulation of (3) aggregates the $2^{J}$ unique response patterns (weighted by their frequencies). We do not adopt this perspective because it becomes less suitable, both computationally and theoretically, for high-dimensional data with very large $J$. On the computational side, our approach does not rely on evaluating the likelihood through these unique response patterns. On the theoretical side, our spectral method is motivated by the approximate low-rank structure of the $N\times J$ binary response matrix $\mathbf{R}\in\mathbb{R}^{N\times J}$ itself, which does not utilize the unique response patterns.
Figure 1: An illustration for spectral clustering: row vectors of $\mathbf{U}\mathbf{\Sigma}$ (left) and $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ (right). Setting: $N=500$, $J=250$, $K=3$.
@@
2Perform rank-$K$ SVD on $\mathbf{R}$ to obtain $\sum_{i=1}^{K}\widehat{\lambda}_{i}\widehat{\bm{u}}_{i}\widehat{\bm{v}}_{i}^{\top}$, where $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq\cdots\geq\widehat{\lambda}_{K}\geq 0$, $\widehat{}\mathbf{U}=\left({\bm{u}}_{1},\cdots,{\bm{u}}_{K}\right)\in\mathbb{R}^{N\times K}$ and $\widehat{}\mathbf{V}=\left({\bm{v}}_{1},\cdots,{\bm{v}}_{K}\right)\in\mathbb{R}^{J\times K}$.
Perform $K$-means on the rows of $\widehat{}\mathbf{U}\widehat{}\mathbf{\Sigma}$ (or equivalently, $\mathbf{R}\widehat{}\mathbf{V}$), i.e.,
@@ 3.2 One-step Likelihood Refinement with Spectral Initialization
    (Assignment) Given $\widehat{}\mathbf{\Theta}$, the objective function $L({\bm{s}},\widehat{}\mathbf{\Theta}\mid\mathbf{R})$ can decoupled into $N$ terms each corresponding to an individual data point, and it is equivalent to assign label $k\in[K]$ to sample $i$ which maximize the log-likelihood independently for each $i\in[N]$:
@@
$\widehat{\theta}_{j,k}=\frac{\sum_{i\in[N]}R_{i,j}\mathbb{I}\left(\widetilde{s}_{i}=k\right)}{\sum_{i\in[N]}\mathbb{I}\left(\widetilde{s}_{i}=k\right)},\qquad j\in[J],~k\in[K].$ (7)
@@
$\widehat{s}_{i}=\operatorname*{arg\,max}_{k\in[K]}\sum_{j\in[J]}\Big[R_{i,j}\log\Big(\widehat{\theta}_{j,k}\Big)+(1-R_{i,j})\log\Big(1-\widehat{\theta}_{j,k}\Big)\Big],\qquad i\in[N].$ (8)
@@
Interestingly, our two-stage procedure resonates with the spirit of the “initialize-and-refine” scheme that has been used in network analysis for stochastic block models (SBMs) (Gao et al., 2017; Lyu et al., 2023; Chen et al., 2022), while the model considered here differs fundamentally in that LCMs handle high-dimensional response data without relying on pairwise interaction structures inherent to SBMs.
In practice, we observe that performing multiple likelihood-based refinement steps may lead to improved convergence and performance compared to the one-step refinement procedure. Thus, we also consider a multiple-step refinement variant of Algorithm 2, which we denote by SOLA+, in the later simulation studies. We set the number of “Maximization ++ Assignment” refinement steps for SOLA+ to be $10$ in all our simulations. In principle, one could also adapt the algorithm by monitoring the change in the joint log-likelihood values and stopping the algorithm once the absolute difference between successive iterations falls below a small threshold (Di Mari et al., 2023).
@@ 3.3 Extension: Classification EM refinement
Our proposed framework can be readily extended to one-step Classification EM (CEM) refinement, where CEM is another widely used approach to estimate latent class models (Celeux and Govaert, 1992; Zeng et al., 2023). It can be regarded is a variant of the standard EM algorithm designed specifically for clustering. Unlike traditional EM, which calculates “soft” assignments by estimating the probability that each observation belongs to every latent class, CEM uses “hard” assignments: each observation is directly assigned to the class with the highest posterior probability. By explicitly modeling the mixture proportion parameters for the latent classes, CEM adjusts the likelihood to account for varying cluster sizes. Concretely, in our setting, one can adapt the one-step refinement in Algorithm 2 to a CEM step by including class proportions. Besides updating $\widehat{\theta}_{j,k}$ as in (7), we also estimate the proportion parameters $\widehat{{\bm{p}}}=(\widehat{p}_{1},\ldots,\widehat{p}_{K})\in[0,1]^{K}$ by
@@
This one‐step CEM update replaces the likelihood‐refinement stage of our two‐stage estimator: start with spectral clustering, then perform the CEM step above. We provide its theoretical guarantee in Section 4. In practice, one could easily generalize the above CEM procedure to a multiple-step refinement variant, which is analogous to SOLA+.